OpenStudy (anonymous):

Find the digit that makes _2,530 divisible by 9 a.6 b.5 c.7 d.8

OpenStudy (popitree):

for a no to be divisible by 9 its individual digits should sum up to a no divisible by 9

OpenStudy (shinalcantara):

you have to do trial and error by replacing the values as indicated by the choices.. anyone knows how to do it not using trial and error?

OpenStudy (amistre64):

since 10^n = 1 (mod 9) then: a10^k + b10^(k-1) + ...+m10^0 = a(1) + b(1) + ...+m(1) (mod 9)

OpenStudy (amistre64):

try and see is the best method at the moment:

OpenStudy (popitree):

say if sum of digits =9 or 18 or 27 etc the num should be divisible by 9

OpenStudy (shinalcantara):

haha yeah ryt